OpenStudy (anonymous):

Assume that p,q, and r are positive integers satisfying the following conditions (1) p>q>r, (2) p,q, and are all primes (3) p-q=r. This information is sufficient to determine the value of which of them? a) none of them b) p only c) q only d) r only e) p,q, and r

OpenStudy (anonymous):

i am confused are p and q primes, or is r a prime also?

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

which is it? is r a prime?

OpenStudy (anonymous):

Yes p,q, and r are all primes

OpenStudy (anonymous):

ok here is a huge hint if p and q are prime numbers, then if they are odd numbers what kind of number do you get when you subtract two odd numbers?

OpenStudy (anonymous):

odd?

OpenStudy (anonymous):

No even

OpenStudy (freckles):

7-5? 21-19? in general (2k+1)-(2i+1)=2k+1-2i-1=2(k-i) k-i is an integer and 2 times any integer gives us an even number

OpenStudy (anonymous):

Ok thank u

OpenStudy (anonymous):

So the answer is D) r only right?

OpenStudy (freckles):

do you know what r is?

OpenStudy (freckles):

Just for fun.

OpenStudy (anonymous):

Nvm then...

OpenStudy (freckles):

Well you are given r is even and r is also prime

OpenStudy (anonymous):

Ok thanks >-<

OpenStudy (freckles):

So that can only make r what?

OpenStudy (anonymous):

prime?

OpenStudy (freckles):

like what number is even and prime

OpenStudy (anonymous):

2

OpenStudy (freckles):

isn't r prime?

OpenStudy (freckles):

yes

OpenStudy (anonymous):

ooh ooh i know pick me!

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

Can u also help me with another problem? I asked my brother and he also said he doesnt know..

OpenStudy (anonymous):

r is prime, r is even, so r is ... p and q could be any twin primes

OpenStudy (anonymous):

even the evil twin primes f and m

OpenStudy (anonymous):

XD Mind helping me with another problem?

OpenStudy (anonymous):

we will both help i am sure

OpenStudy (anonymous):

:D